# Reclaiming “Open AI" - AI Model Serving Can Be Open Access, Yet Monetizable and Loyal

### v3

@@ 4. $\star$
The OML 1.0 system operates on crypto-economic principles: model users (e.g., services hosting the model publicly) agree to usage terms and post collateral. Authorized provers can then query these hosted models with the secret keys. If a model user provides a fingerprinted response without logged authorized usage (e.g., permission from an access layer), a violation is proven. This can lead to collateral slashing, deterring unauthorized use or distribution, as each distributed model copy contains unique fingerprints. OML 1.0 prioritizes efficiency and offers "next-day security," where compliance is incentivized through the high likelihood of post-violation detection and significant penalties, inspired by optimistic security principles. Further details on OML 1.0’s workflow, specific implementation, and experimental results are elaborated in Appendix B.
Beyond individual model protection with OML primitives like OML 1.0, fostering an incentive-aligned ecosystem requires a broader framework. We conceptually outline a multi-layered, blockchain-based end-to-end protocol designed for this purpose, and a detailed exploration of this end-to-end framework, including its workflow and security analysis, is available in Appendix C.
4 Alternative Views
In this section, we explore alternative perspectives that question the necessity and feasibility of OML, and we articulate our vision by critically engaging with these viewpoints.
4.1 Necessity: Are Existing Approaches Sufficient?
Some may argue OML is unnecessary, as existing paradigms may already form an efficient market equilibrium. OpenAI’s black-box API model (1.2) offers convenience, rapid iteration, and straightforward monetization, appealing to users prioritizing ease and performance. Conversely, HuggingFace’s fully open-weight approach (1.2) champions transparency, reproducibility, and community-driven innovation, allowing extensive local customization. These distinct models currently cater to different, significant segments of the AI landscape, raising questions on the need of a new framework.
4.2 Feasibility: Economic and Technical Concerns
Feasibility concerns also challenge OML. Theoretical OML constructions discussed in Section 3.1, such as those using advanced cryptography or TEEs, can introduce significant cost and latency, potentially hindering democratized AI access. More practical solutions like OML 1.0, while reducing overheads via fingerprinting, offer weaker, post-hoc security guarantees, and may not prevent sophisticated attacks. Furthermore, OML does not fully eliminate risks from malicious model creators embedding hidden backdoors prior to model release to create biased or downgraded responses. Finally, some may argue that policy-based alternative, like regulatory audits and licensing, could more effectively address misuse and ethical concerns without technical overheads inherent to OML.
4.3 Our Vision
Acknowledging the distinct strengths of current paradigms: closed-source APIs (e.g., OpenAI), fully open-weight platforms (e.g., Hugging Face), and policy-driven frameworks, our vision for OML engages these as follows:
  • While closed systems and policy approaches have roles, over-reliance on dominant corporations or centralized mandates risks fragile, distorted AI futures. OML counters by promoting an idealized ecosystem with open accessibility and robust monetizability, fostering collective management and sustainable rewards to prevent singular dominance.
  • While the open-weight paradigm fuels collaborative innovation, it often struggles with sustainable funding for foundational model creation. OML aims for economic self-sufficiency, empowering foundational model builders, not just application developers, to continuously produce and refine high-quality models.
We recognize significant open questions regarding technical limitations and creator-related vulnerabilities, such as backdoor risks (pertinent to many open-weight models), the substantial challenges in balancing cost, performance, and security for OML constructions, and the hardness of finding and embedding robust and undetectable fingerprints into AI models. Precisely due to these complexities, this paper serves as a Call for Extensive Research into more robust, efficient, user-friendly, and secure OML designs. By outlining OML’s conceptual framework, we aim to stimulate broader inquiry into its theory, feasibility, and viability. Our aspiration is for OML to cornerstone a fair, resilient, self-sustaining AI ecosystem, paving the way for equitable benefits.
@@ 5 Conclusion
In this position paper, we introduce the Open-access, Monetizable, and Loyal (OML) paradigm as a framework crucial for the fair distribution, sustainable deployment, and equitable governance of AI models. The paper underscores the necessity for OML (Section 1), formalizes its core properties and definitions (Section 2), and explores theoretical constructions alongside practical pathways, including OML 1.0 and a supporting ecosystem (Sections 3). By proposing OML, we aim to catalyze extensive research and collaborative initiatives towards developing next-generation AI systems that are open, resilient, and foster a truly equitable and beneficial AI future for all.
Appendix A Canonical OML Constructions
A.1 Obfuscation
@@ A.2 Fingerprinting
Our first practical strategy, which we call OML 1.0, builds upon this fingerprinting technique, which we introduce in Appendix B.
@@
An elaborate version of this approach is presented in Appendix B as OML 1.0.
A.3 Trusted Execution Environments (TEEs)
@@
OML formatting. As before, we can use any cryptographic scheme $(Enc_{pk},Dec_{sk})$ where the secret key $sk$ is only accessible within the TEE. The model is wrapped in a program that executes the desired task (e.g. inference or fine-tuning) conditioned on the $Verify_{sk}$ function as usual. This program is then encrypted with a public key before it is published onto the Sentient protocol as a TEE-based OML format. After a user is granted access to download this OML file by the Sentient protocol, the user can launch the TEE application using the Sentient SDK on any TEE-enabled machine. The SDK manages the launch of the program with a decryptor module inside a TEE. The SDK and the decryptor module coordinate the secure transfer of the private key directly into the unaltered TEE runtime to decrypt the model inside the TEE.
Verification and Usage. First, the user requests a permission string $\sigma$ from the Sentient protocol by sending $h(x)$ to it for some input $x$. Afterwards, the user can pass $(x,\sigma)$ into the TEE via a secure channel by using the Sentient SDK. The OML file inside the TEE will then verify the permission string $\sigma$, run the task on input $x$ and provide the result back to the user.
Additional requirement. This OML implementation must provide a guarantee that the program running inside the TEE is unmodified by a malicious user. This is to ensure that any and all data or intermediate results during the execution of the .$oml$ file inside the secure program are not retrievable by a malicious user. More precisely, the secure program must be exactly the program that was constructed by an honest SDK from the published OML file. Whether or not the process has been modified can be verified by the hash of the program with remote attestation.
@@ A.4 Cryptography
Verification and Usage. In FHE and HE, for an inference request from the user with input $x$, users first request the permission $\sigma(x)=Enc_{k1}(x)$ from Sentient, then run inference with the OML file on encrypted data $\sigma(x)$, and finally send the final result to the Sentient platform for decryption to plain-text results. In FE, users first request the permission string $\sigma(x)=Enc_{k1}(x)$, and then locally run the OML file on the permission string $\sigma(x)$ to get the desired output.
@@
A.5 Melange – an OML Construction with a Mixture of Security Guarantees
@@ A.5.1 Example Workflow
Verification and Usage phase. For a user who wants to do an inference task, we follow the methods from Sections A.3 and A.4 to locally run the inference task (and thus protected by cryptographic or hardware guarantees); we execute the obfuscated binary file for inference of the layers in $M_{2}$ (and thus inherit obfuscation guarantees described in Section A.1).
A.5.2 Security Analysis
@@ A.6 Summary
[…]
@@
Appendix B OML 1.0: Turning Attack Methods on AI into a Security Tool
@@
In this chapter we expand upon the optimistic version of OML (introduced in Section A.2) as OML 1.0. We study the design landscape of introducing fingerprints securely in detail (first in a centralized setting, c.f. Section B.1). We conduct a detailed security analysis of OML 1.0, paying close attention to coalition attacks; we show in Section B.2 that an adaptive fingerprint querying scheme in OML 1.0 makes it secure against this formidable attack vector. We generalize the OML 1.0 approach to a decentralized scenario in Section B.3.
B.1 Sentient Protocol under a Single Trusted Prover

### v4

@@ 3.1 Canonical Constructions: Security-Performance Spectrum
Detection economics. If a host fails to authorize an $\alpha$ fraction of public queries, the probability of evading $n$ independent fingerprints is $\Pr[\text{undetected}]=(1-\alpha)^{n},\quad\text{so}\quad\Pr[\text{caught}]=1-(1-\alpha)^{n}.$ Choosing $n$ and the probing cadence to make $(1-\alpha)^{n}\!\ll\!10^{-k}$ yields predictable compliance guarantees against rational adversaries, with zero runtime overhead on ordinary traffic.
OMLization recipe. We adopt a two-objective fine-tuning: (i) bind $k_{i}\!\mapsto\!r_{i}$; (ii) minimize task loss on $\mathcal{D}_{\mathcal{X}}$ with anti-forgetting regularizers (e.g., rehearsal, weight-averaging). To harden against serving stacks, we prompt-augment fingerprints using system-prompt templates expected in deployment (role prompts), and place keys both in-distribution (stealth) and slightly out-of-distribution (capacity).
Algorithm 5 OMLize-Fingerprint (OML 1.0): training and enforcement
1:Input: base model $M$, secret $\mathcal{K}_{\mathrm{fp}}=\{(k_{i},r_{i})\}_{i=1}^{n}$, task data $\mathcal{D}$, anti-forgetting params
2:Training loop: minimize $\mathcal{L}=\lambda_{\mathrm{task}}\mathcal{L}_{\mathrm{task}}(M;\mathcal{D})+\lambda_{\mathrm{fp}}\tfrac{1}{n}\sum_{i}\ell\big(M(k_{i}),r_{i}\big)+\lambda_{\mathrm{af}}\mathcal{R}_{\mathrm{anti\text{-}forget}}$
3:Prompt augmentation: sample serving templates $\pi$ and train on $\pi(k_{i})\mapsto r_{i}$ for robustness
4:Platform: issue per-input tokens, log authorized uses (commitments to $h(x)$), escrow collateral
5:Prover cadence: probe a random subset of $\mathcal{K}_{\mathrm{fp}}$; slash collateral on verified violations
6:Publish: release $M_{\mathrm{oml}}$ (weights) + policy; keep $\mathcal{K}_{\mathrm{fp}}$ secret
We summarize the key experimental findings that show the feasibility of OML 1.0. Full setups, experiment details, ablations, security analysis, and plots can be found in Appendix C.
Utility vs. capacity. A central question is how many fingerprints can be embedded before harming task performance. On a 7B-scale base model evaluated on standard language tasks, we observe that up to $\sim\!10^{3}$ fingerprints can be embedded with near-baseline accuracy when using anti-forgetting and weight-averaging. This gives an initial operating region for $n$ that balances detection and accuracy.
Persistence under benign fine-tuning. Hosts often fine-tune for their domain. We fine-tune post-OMLization models on a standard SFT corpus and measure fingerprint survival and utility. A substantial fraction of fingerprints persists (e.g., $\gtrsim\!50\%$ at $\leq\!2$K capacity), and downstream utility remains within a few points of baseline, supporting the post-hoc accountability.
Together, these results show that OML 1.0 is deployable today with post-hoc enforcement: it preserves utility at useful capacities, remains robust to realistic serving perturbations and benign fine-tuning, and provides tunable, high-confidence enforcement with negligible inference overhead.
Deployment Synthesis. Table 2 summarizes the complete construction spectrum and their tradeoffs. The canonical constructions provide a progressive hardening path for how OML gets realized: begin with OML 1.0 for immediate needs, identify critical components through usage analysis, and selectively apply stronger protections as infrastructure matures and threats evolve.
Table 2: OML construction summary. Symbols: $\checkmark~\text{strong},\triangle~\text{partial},\circ~\text{low}$.
Construction Control White-box Robust Overhead Readiness Core Assumption
Obfuscation Pre-hoc $\triangle$ $\circ$ Negligible Immediate Security by obscurity
TEE-gated Pre-hoc $\checkmark$ $\checkmark$ Moderate Rising Hardware trust
Cryptographic Pre-hoc $\checkmark$ $\checkmark$ Very High Limited No extra trust
Melange Pre-hoc $\checkmark$ $\triangle$ $\triangle$ Immediate Component union
OML 1.0 Post-hoc $\triangle$ Low Immediate Ecomomic deterrent
A more detailed analysis on canonical OML constructions can be found in Appendix B, while extensive experiments alongside OML 1.0 and security analysis can be found in Appendix C.
@@ 4 Conclusion
In this paper, we introduce and formalize the OML primitive as a foundation for fair distribution, sustainable deployment, and accountable governance of AI models. We articulate its core properties, establish theoretical limits and sufficient conditions, and outline a practical path with empirical evidence. Our goal is to crystallize OML as a coherent research direction with significant implications for how AI capabilities are distributed, monetized, and governed.
Appendix A Proofs for Section 2.3
This appendix provides proofs for Theorems 2.22.4. We begin by stating the technical assumptions used in Sec. 2.3 and here.
Standing assumptions.
Unless otherwise specified, we consider models $M:\mathcal{X}\to\mathcal{Y}$ with $\mathcal{X}\subseteq\mathbb{R}^{d}$ measurable and $\mathcal{Y}\subseteq[0,1]^{m}$ bounded. The performance metric $d:\mathcal{Y}\times\mathcal{Y}\to[0,1]$ is either: (i) a coordinatewise squared loss with averaging, i.e. $d(u,v)=\tfrac{1}{m}\sum_{j=1}^{m}(u_{j}-v_{j})^{2}$, or (ii) a bounded Bregman divergence $D_{\phi}(u,v)$ induced by a $1$-strongly convex, $L$-smooth $\phi$ on a compact convex subset of $0,1]^{m}$ (so $D_{\phi}\in[0,1]$ after normalization). For classification we also consider 011 loss. Hypothesis classes $\mathcal{H}\subseteq[0,1]^{m}{}^{\mathcal{X}}$ have finite pseudo-dimension (real-valued case) or VC dimension (classification). Expectations $\mathbb{E}[\cdot]$ are over $x\sim\mathcal{D}_{\mathcal{X}}$ unless stated. We use standard uniform-convergence and Rademacher-complexity results for bounded function classes [90, 65, 10].
A.1 Proof of Theorem 2.2 (Information-theoretic impossibility)
Proof A.1.
Finite domain. If $|\mathcal{X}|<\infty$, an unbounded adversary with unlimited authorization access enumerates all $x\in\mathcal{X}$, obtains $p_{x}=\sigma(h(x),k_{\text{own}})$, evaluates $y_{x}=M_{oml}(x,p_{x})=M(x)$, and stores $T(x)=y_{x}$. The table $T$ replicates $M$ exactly thereafter, without authorization.
General domain. Suppose $\mathcal{X}\subseteq\mathbb{R}^{d}$, $\mathcal{Y}\subseteq[0,1]^{m}$, and $M$ is measurable. With unbounded queries, the adversary draws i.i.d. $x_{i}\sim\mathcal{D}_{\mathcal{X}}$, obtains $y_{i}=M(x_{i})$ from authorized evaluations, and fits $\hat{h}$ by empirical risk minimization over a hypothesis class $\mathcal{H}$ containing $M$ (e.g., the realized architecture family). For squared loss or a bounded Bregman divergence, uniform convergence yields for some constant $C>0$:
$\sup_{h\in\mathcal{H}}\Big|\ \mathbb{E}\,d\big(h(x),M(x)\big)-\tfrac{1}{N}\!\sum_{i=1}^{N}d\big(h(x_{i}),M(x_{i})\big)\ \Big|\ \leq\ C\sqrt{\frac{\mathrm{Pdim}(\mathcal{H})+\log(1/\delta)}{N}}$
with probability $\geq 1-\delta$ [65, 90, Chs. 3,11]. As $N\to\infty$, the right-hand side vanishes, and ERM (or structural risk minimization) produces $\hat{h}$ with $\mathbb{E}\,d(\hat{h}(x),M(x))\to 0$. Thus, information-theoretic protection is impossible against an unbounded adversary with unlimited oracle access.
A.2 Proof of Theorem 2.3 (OML from indistinguishability obfuscation)
Proof A.2 (Construction and argument).
Assume an indistinguishability obfuscation (iO) scheme for the relevant circuit class and a signature scheme $\mathsf{Sig}=(\mathsf{KeyGen},\mathsf{Sign},\mathsf{Verify})$ that is EUF-CMA secure. Define the circuit
$C(x,p)\;=\;\begin{cases}M(x)&\text{if }\ \mathsf{Verify}\big(h(x),p,vk_{\text{own}}\big)=1,\\[2.0pt] \mathsf{Noise}(x)&\text{otherwise},\end{cases}$
where $\mathsf{Noise}$ is any efficiently computable low-utility mapping whose range lies in a small $d$-ball around a baseline (e.g., a fixed vector, or a pseudorandom output independent of $x$). Publish $M_{oml}\triangleq iO(C)$ and $vk_{\text{own}}$; retain $k_{\text{own}}$.
Forgery resistance. Suppose a PPT adversary produces $(x^{*},p^{*})$ without prior authorization for $x^{*}$ such that $d\big(M_{oml}(x^{*},p^{*}),M(x^{*})\big)\leq\epsilon_{\text{utility}}$. Functionality preservation under iO implies $\mathsf{Verify}(h(x^{*}),p^{*},vk_{\text{own}})=1$, yielding an existential forgery for $\mathsf{Sig}$ on message $h(x^{*})$, contradicting EUF-CMA.
Extraction resistance. Let $C_{0}$ be the circuit above and $C_{1}$ be any syntactically distinct circuit computing the same function (e.g., with inlining and control-flow reorganization that entangles verification with model computation). By iO, $iO(C_{0})$ and $iO(C_{1})$ are computationally indistinguishable. Any white-box procedure that reliably identifies and removes verification logic from $iO(C_{0})$—thereby recovering a high-utility version of $M$ that bypasses authorization—would also work on $iO(C_{1})$, where such separation is obfuscated by construction, contradicting indistinguishability. Hence, under iO and EUF-CMA, a PPT adversary cannot (i) forge tokens to obtain authorized outputs on fresh inputs nor (ii) recover a functionally equivalent, authorization-free model producing high-utility outputs. This establishes the stated guarantees.
Remark.
The signature scheme prevents functional bypass; iO prevents structural separability of verification under white-box access. Practical OML designs approximate these guarantees via verifier entanglement, TEEs, FHE/MPC hybrids, or melange constructions; iO is used here as a sufficiency anchor, not as a practical prescription.
A.3 Proof of Theorem 2.4 (Query–security trade-off)
We treat the real-valued case under squared loss; the bounded Bregman case follows by identical symmetrization (boundedness ensures the same $1/\sqrt{N}$ rate up to constants), and the realizable 0–1 classification case yields the standard $1/\varepsilon$ dependence.
Lemma A.3 (Uniform convergence under squared loss).
Let $\mathcal{H}\subseteq[0,1]^{m}{}^{\mathcal{X}}$ with pseudo-dimension $d$. There exists $C_{1}>0$ such that, for any $\delta\in(0,1)$ and i.i.d. sample $(x_{i})_{i=1}^{N}$,
$\Pr\Bigg[\ \sup_{h\in\mathcal{H}}\Big|\,\mathbb{E}\ d\big(h(x),M(x)\big)-\frac{1}{N}\sum_{i=1}^{N}d\big(h(x_{i}),M(x_{i})\big)\,\Big|\ \leq\ C_{1}\sqrt{\frac{d+\log(1/\delta)}{N}}\ \Bigg]\ \geq\ 1-\delta.$
Proof A.4.
Let $\mathcal{L}=\{\ell_{h}(x)=d(h(x),M(x)):h\in\mathcal{H}\}\subseteq[0,1]$. The VC-subgraph dimension of $\mathcal{L}$ is $O(d)$ (closure of pseudo-dimension under Lipschitz maps on bounded ranges). Standard VC-subgraph or Rademacher complexity bounds yield the inequality; see [65, Ch. 3] and [90, Ch. 11].
Proof A.5 (Proof of Theorem 2.4).
Assume realizability: $M\in\mathcal{H}$. Let the adversary observe $N$ authorized pairs $(x_{i},y_{i})$ with $y_{i}=M(x_{i})$ and return an empirical risk minimizer
$\hat{h}\in\arg\min_{h\in\mathcal{H}}\ \frac{1}{N}\sum_{i=1}^{N}d\big(h(x_{i}),M(x_{i})\big).$
By realizability, the empirical risk of $M$ is 0, so the empirical risk of $\hat{h}$ is $\leq 0$. Applying Lemma A.3 to $\hat{h}$ and using nonnegativity of $d$,
$\mathbb{E}\ d\big(\hat{h}(x),M(x)\big)\ \leq\ C_{1}\sqrt{\frac{d+\log(1/\delta)}{N}}$
with probability at least $1-\delta$. Therefore, $\mathbb{E}\,d(\hat{h}(x),M(x))\leq\varepsilon$ whenever $N\geq C\,\frac{d+\log(1/\delta)}{\varepsilon^{2}}$ with $C=C_{1}^{2}$, proving the positive direction. The operational converse follows by contrapositive: to preclude $(\varepsilon,\delta)$-accurate extraction by such ERM adversaries, an OML deployment must enforce $N<c\,\frac{d+\log(1/\delta)}{\varepsilon^{2}}$ for some absolute $c>0$ (absorbing constants).
Remark.
(1) Bounded Bregman divergences. If $d=D_{\phi}$ with $\phi$ 11-strongly convex and $L$-smooth on a compact convex domain and $D_{\phi}\in[0,1]$ (after normalization), the same uniform convergence rate holds using the Lipschitzness of $\ell_{h}(x)=D_{\phi}(h(x),M(x))$ in its arguments. (2) 0–1 loss. In realizable binary classification with VC dimension $d$, the optimal sample complexity is $\Theta((d\log(1/\varepsilon)+\log(1/\delta))/\varepsilon)$ [90]; the OML constraint is analogous with the $1/\varepsilon$ dependence.
A.4 Additional technical notes
On $\mathsf{Noise}$.
Any fixed choice with low expected utility (e.g., constant output or PRG-based mapping independent of $x$) suffices; boundedness ensures compatibility with the metric normalization in the main text.
On public parameters.
Exposing $vk_{\text{own}}$ enables decentralized verification of authorization. The proofs above do not require $vk_{\text{own}}$ to be hidden; secrecy resides solely in $k_{\text{own}}$.
On realizability.
The trade-off in Theorem 2.4 is stated under realizability to isolate the effect of authorized answers. Under agnostic noise, replace ERM bounds with excess risk bounds; the qualitative inverse-square dependence on $\varepsilon$ for bounded real-valued losses remains (up to constants).
Appendix B Canonical OML Constructions
B.1 Obfuscation
@@ B.2 Fingerprinting
Our first practical strategy, which we call OML 1.0, builds upon this fingerprinting technique, which we introduce in Appendix C.
@@
An elaborate version of this approach is presented in Appendix C as OML 1.0.
B.3 Trusted Execution Environments (TEEs)
@@
OML formatting. As before, we can use any cryptographic scheme $(Enc_{pk},Dec_{sk})$ where the secret key $sk$ is only accessible within the TEE. The model is wrapped in a program that executes the desired task (e.g. inference or fine-tuning) conditioned on the $Verify_{sk}$ function as usual. This program is then encrypted with a public key before it is published onto the auditing protocol as a TEE-based OML format. After a user is granted access to download this OML file by the auditing protocol, the user can launch the TEE application on any TEE-enabled machine. The SDK manages the launch of the program with a decryptor module inside a TEE. The SDK and the decryptor module coordinate the secure transfer of the private key directly into the unaltered TEE runtime to decrypt the model inside the TEE.
Verification and Usage. First, the user requests a permission string $\sigma$ from the auditing protocol by sending $h(x)$ to it for some input $x$. Afterwards, the user can pass $(x,\sigma)$ into the TEE via a secure channel by using the SDK. The OML file inside the TEE will then verify the permission string $\sigma$, run the task on input $x$ and provide the result back to the user.
Additional requirement. This OML implementation must provide a guarantee that the program running inside the TEE is unmodified by a malicious user. This is to ensure that any and all data or intermediate results during the execution of the .oml.oml file inside the secure program are not retrievable by a malicious user. More precisely, the secure program must be exactly the program that was constructed by an honest SDK from the published OML file. Whether or not the process has been modified can be verified by the hash of the program with remote attestation.
@@ B.4 Cryptography
Verification and Usage. In FHE and HE, for an inference request from the user with input $x$, users first request the permission $\sigma(x)=Enc_{k1}(x)$ from the platform, then run inference with the OML file on encrypted data $\sigma(x)$, and finally send the final result to the auditing platform for decryption to plain-text results. In FE, users first request the permission string $\sigma(x)=Enc_{k1}(x)$, and then locally run the OML file on the permission string $\sigma(x)$ to get the desired output.
@@
B.5 Melange – an OML Construction with a Mixture of Security Guarantees
@@ B.5.1 Example Workflow
Verification and Usage phase. For a user who wants to do an inference task, we follow the methods from Sections B.3 and B.4 to locally run the inference task (and thus protected by cryptographic or hardware guarantees); we execute the obfuscated binary file for inference of the layers in $M_{2}$ (and thus inherit obfuscation guarantees described in Section B.1).
B.5.2 Security Analysis
@@ B.6 Summary
Obfuscation $\mathbf{\times}$ $\mathbf{\times}$
Fingerprinting ✓($\mathbf{\times}$ if monetizable)
TEEs $\mathbf{\times}$ ✓or $\mathbf{\times}$ $\mathbf{\times}$
Cryptography $\mathbf{\times}$ $\mathbf{\times}$
@@
Appendix C OML 1.0: Turning Attack Methods on AI into a Security Tool
@@
In this chapter we expand upon the optimistic version of OML (introduced in Section B.2) as OML 1.0. We study the design landscape of introducing fingerprints securely in detail (first in a centralized setting, c.f. Section C.1). We conduct a detailed security analysis of OML 1.0, paying close attention to coalition attacks; we show in Section C.2 that an adaptive fingerprint querying scheme in OML 1.0 makes it secure against this formidable attack vector. We generalize the OML 1.0 approach to a decentralized scenario in Section C.3.
C.1 The Auditing Protocol under a Single Trusted Prover